In order to at least break-even, a price must cover total costs (COGS + fixed costs). Calculate total cost and per unit total cost for XYZ eyeglass manufacturer with the following information (show your calculations):

·      Total units produced and sold: 100,000

·      Total annual management compensation:       $600,000

·      Annual utilities cost:                        $120,000

·      Annual office rent:                        $240,000

·      Annual advertising cost:                  $400,000

·      Annual shared admin department costs:      $600,000

·      20,000 plastic frames purchased in March for $9 each

·      40,000 plastic frames purchased in August for $11 each

·      100,000 units of screws purchased in January for $1 per unit
Answer #1

Total fixed annual cost = 600000+120000+240000+400000+600000

= $ 1,960,000

Total COGS = 20000*9+40000*11+100000*1= $ 720,000

Total fixed annual cost + COGS = 1960000+720000

= $ 2,680,000

Break-even price = (Total fixed annual cost + COGS)/Number of units sold

= 2680000/100000

= $ 26.80